# Class-Conditional Distribution Balancing for Group Robust Classification

### v2

@@ Abstract
Spurious correlations that lead models to correct predictions for the wrong reasons pose a critical challenge for robust real-world generalization. Existing research attributes this issue to group imbalance and addresses it by maximizing group-balanced or worst-group accuracy, which heavily relies on expensive bias annotations. A compromise approach involves predicting bias information using extensively pretrained foundation models, which requires large-scale data and becomes impractical for resource-limited rare domains. To address these challenges, we offer a novel perspective by reframing the spurious correlations as imbalances or mismatches in class-conditional distributions, and propose a simple yet effective robust learning method that eliminates the need for both bias annotations and predictions. With the goal of reducing the mutual information between spurious factors and label information, our method leverages a sample reweighting strategy to achieve class-conditional distribution balancing, which automatically highlights minority groups and classes, effectively dismantling spurious correlations and producing a debiased data distribution for classification. Extensive experiments and analysis demonstrate that our approach consistently delivers state-of-the-art performance, rivaling methods that rely on bias supervision.
Index Terms:
Spurious correlations, group robust learning, sample reweighting, distribution balancing, stable learning.
I Introduction
@@
Despite the unprecedented success of deep learning in recent decades, machine learning methods following the ERM paradigm [1] often suffer from spurious correlations when training data is biased [2, 3]: some spurious factors (also termed bias) are predictive of the target class solely due to their high co-occurrence with the target label, despite being irrelevant to the true classification function. It causes unpredictable performance losses under distribution shift, where spurious correlations in the test data differ from or conflict with those in the training data [4, 5, 6]. For example, in the task of recognizing cows and camels [7, 8], the training dataset is biased with camel images predominantly associating with desert backgrounds, leaving camels on grass as a minority group [9]. Consequently, the classifier is likely to associate camels with deserts, resulting in a biased model that fails the underrepresented groups [10, 11, 12]. Beyond such straightforward scenarios, spurious correlations have emerged as a significant concern in several high-stakes domains, including visual question answering [13], autonomous driving [14], facial recognition [15], criminal justice systems [16], financial risk modeling [17], and healthcare diagnostics [18], where the imperative to ensure fairness is especially pronounced [19, 20].
@@
While progress has been made, existing research attributes spurious correlations to group imbalance. They enhance robustness against spurious correlations by utilizing bias information to implement group balanced measures during training [21, 22, 23] or validation [24, 25, 26, 27, 28]. These methods can be categorized into bias-supervised and pseudo-bias-supervised methods, depending on how bias information is obtained. Bias-supervised methods [21, 22, 24, 25, 26, 27, 28] rely on ground-truth bias annotations provided by human annotators, which limits their applicability in practical scenarios where such annotations are too costly or even impossible to obtain. To address these limitations, pseudo-bias-supervised methods [29, 30, 2, 23, 8, 31, 32] are developed to use predicted bias information instead of human annotations. By leveraging extensively pretrained base models [33] or large scale visual language models [34, 35], some methods achieve performance only slightly inferior to that of bias-supervised approaches. However, the pretraining process is both data- and resource-intensive, making it impractical for rare domains(e.g., hyperspectral images [36], medical images [37], and point cloud data [38]), where collecting large-scale paired vision-language datasets is impractical.
@@
Here we consider a more realistic and challenging scenario where neither bias annotations nor a pretrained model for bias prediction are available. With this goal, a line of research on stable learning has been proposed [5, 39, 40, 41, 42]. They make a structural assumption about data features, splitting them into causal and spurious factors, and hypothesize that the statistical dependence between causal and spurious factors is the primary cause of model failure under distribution shift. These methods propose a global sample reweighting strategy to remove spurious correlations by decorrelating all features, without requiring explicit bias information. However, with finite samples, achieving perfect independence is in theory challenging. Moreover, most of these methods are designed for linear frameworks with predefined features. When extended to deep model, however, their effectiveness diminishes due to the heavy entanglement of learned features, where mere decorrelation fails to ensure proper disentanglement [40].

### v3

@@ Abstract
Spurious correlations that lead models to correct predictions for the wrong reasons pose a critical challenge for robust real-world generalization. Existing research attributes this issue to group imbalance and addresses it by maximizing group-balanced or worst-group accuracy, which heavily relies on expensive bias annotations. A compromise approach involves predicting bias information using extensively pretrained foundation models, which requires large-scale data and becomes impractical for resource-limited rare domains. To address these challenges, we offer a novel perspective by reframing the spurious correlations as imbalances or mismatches in class-conditional distributions, and propose a simple yet effective robust learning method that eliminates the need for both bias annotations and predictions. With the goal of maximizing the conditional entropy (uncertainty) of the label given spurious factors, our method leverages a sample reweighting strategy to achieve class-conditional distribution balancing, which automatically highlights minority groups and classes, effectively dismantling spurious correlations and producing a debiased data distribution for classification. Extensive experiments and analysis demonstrate that our approach consistently delivers state-of-the-art performance, rivaling methods that rely on bias supervision.
1 Introduction
@@
Despite the unprecedented success of deep learning in recent decades, machine learning methods following the ERM paradigm [1] often suffer from spurious correlations when training data is biased: some spurious factors (also termed bias) are predictive of the target class solely due to their high co-occurrence with the target label, despite being irrelevant to the true classification function [2, 3]. It causes unpredictable performance losses under distribution shift, where spurious correlations in the test data differ from or conflict with those in the training data [4, 5, 6]. For example, in the task of recognizing cows and camels [7], the training dataset is biased with camel images predominantly associating with desert backgrounds, leaving camels on grass as a minority group [8]. Consequently, the classifier is likely to associate camels with deserts, resulting in a biased model that fails the underrepresented groups [9]. Moving beyond these simple instances, spurious correlations have surfaced as a major challenge in numerous high-stakes settings. This issue is especially evident in areas such as criminal justice systems [10], autonomous driving [11], and medical diagnosis [12], where ensuring fairness is of paramount importance [13].
@@
While progress has been made, existing research attributes spurious correlations to group imbalance. They enhance robustness against spurious correlations by utilizing bias information to implement group balanced measures during training [14, 15, 16] or validation [17, 18, 19, 20, 21]. These methods can be categorized into bias-supervised and pseudo-bias-supervised methods, depending on how bias information is obtained. Bias-supervised methods [14, 15, 17, 18, 19, 20, 21] rely on ground-truth bias annotations provided by human annotators, which limits their applicability in practical scenarios where such annotations are too costly or even impossible to obtain. To address these limitations, pseudo-bias-supervised methods [22, 23, 2, 16, 7] are developed to use predicted bias information instead of human annotations. By leveraging extensively pretrained base models [24] or large scale visual language models [25, 26], some methods achieve performance only slightly inferior to that of bias-supervised approaches. However, the pretraining process is both data- and resource-intensive, making it impractical for rare domains(e.g., hyperspectral images [27], medical images [28], and point cloud data [29]), where collecting large-scale paired vision-language datasets is impractical.
@@
Here we consider a more realistic and challenging scenario where neither bias annotations nor a pretrained model for bias prediction are available. With this goal, a line of research on stable learning has been proposed [5, 30, 31]. They make a structural assumption about data features, splitting them into core and spurious factors, and hypothesize that the statistical dependence between core and spurious factors is the primary cause of model failure under distribution shift. These methods propose a global sample reweighting strategy to remove spurious correlations by decorrelating all features, without requiring explicit bias information. However, with finite samples, achieving perfect independence is in theory challenging. Moreover, most of these methods are designed for linear frameworks with predefined features. When extended to deep models, however, their effectiveness diminishes due to the heavy entanglement of learned features, where mere decorrelation fails to ensure proper disentanglement [30].